# Bose-Einstein condensation in exotic lattice geometries

### v1

@@ III.2 Critical temperature
In the previous discussion, we have already noted substantial differences of the critical temperature in different geometries at fixed system sizes. From a theoretical point of view, however, the most interesting aspect is the scaling of the critical temperature with the system size. It is well established that, in the absence of a trapping potential, the critical temperature of a 2D systems vanishes in the thermodynamic limit, and remains finite only in 3 (or more) spatial dimensions. In finite lattice systems, this reflects in a behavior where the critical temperature monotonically decays with system size in the square lattice (Fig. 4 (c)), but tends towards a constant value in the cubic lattice (Fig. 4 (b)).
We also observe a monotonous decay of $T_{\rm c}$ vs. system size in all fractal lattices. A direct comparison of the fractal lattices and their regular counterparts is shown in Fig. 4 (a,c). This behavior suggests that for fractal geometries the critical temperature also vanishes in the thermodynamic limit. Obviously, this is not at all surprising for the Sierpiński gasket and carpet, which are embedded in a 2D space and have a Hausdorff dimension below 2. However, we observe a similar decay of critical temperature also in the case of a Sierpiński tetrahedron. This structure is embedded in 3D, and the regular lattice counterpart is characterized by a finite $T_{\rm c}$ in the thermodynamic limit. On the other hand, the Hausdorff dimension of the Sierpiński tetrahedron is exactly 2, and it appears that the critical temperature behaves qualitatively the same as in the other 2D systems.
@@
Strikingly different is the behavior in hyperbolic lattices. Although these lattices can be mapped onto a 2D Poincaré disk, the critical temperatures are found to increase with the system size, for sufficiently large systems, see Fig. 5. Clearly, this behavior suggests that the critical temperature does not vanish in the thermodynamic limit. Instead, we expect that the critical temperature will saturate as in the cubic lattice, however at system sizes that are too large for our computations. In this sense, the hyperbolic lattices provide an remarkable exception to other 2D structures with vanishing $T_{\rm c}$.
@@
Figure 6: Integrated density of states (IDoS) in (a) three fractal lattices: Sierpiński carpet with $V=4096$ sites (black), Sierpiński triangle with $V=3282$ sites (purple), and Sierpiński tetrahedron with $V=2050$ sites (orange), as well as (b) hyperbolic and regular rectangular lattices: cubic lattice with $V=18^{3}=5832$ sites (black), rectangular lattice with $V=80^{2}=6400$ sites (purple), $\{p,q\}=\{3,7\}$ hyperbolic lattice with $V=6615$ sites (orange) and $\{p,q\}=\{7,3\}$ hyperbolic lattices with $V=5887$ sites (blue). For each geometry, the energy of each state $E$ is normalized using the energy of the ground state $\epsilon_{0}$ and the highest energy $\epsilon^{*}=\max(\epsilon_{k})$, so that $\tilde{E}=(E-\epsilon_{0})/(\epsilon^{*}-\epsilon_{0})$. In fractal lattices, the IDoS increases rapidly near the ground state, as shown in the red frame of Fig. 6(a). As can be seen in Fig. 6(b), the IDoS in the rectangular lattice scales linearly near the ground state, while in the cubic lattice, as well as the hyperbolic lattices, it scales quadratically.
@@ III.3 Density of states
The behavior discussed so far is a consequence of the spectral properties in the different geometries. Therefore, let us now have a look at the integrated density of states (IDoS) in the different geometries. In Fig. 6(a) for the fractal lattices, we see that the triangle and tetrahedron share similar properties of the energy spectrum, including large gaps and states clustered together in narrow energy intervals. On the other hand, the Sierpiński carpet has an IDoS more similar to that of a standard square lattice, cf. Fig 6(b). It is not as smooth and thus results in a drastically different density of states. The spectral properties are investigated in more detail by Yao et al. in Ref. [129].
It is worth noting how the IDoS of the 3D cubic lattice behaves very differently from all fractal lattices, see Fig. 6(b). In particular, the quadratic behavior near the ground state distinguishes the IDoS of the cubic lattice from the fractal and the 2D rectangular lattice, displaying a linear behavior. As opposed to quadratic IDoS, a linear IDoS results in a non-zero density of states (DoS) near the ground state, which prevents BEC in the thermodynamic limit [130, 131].
From this perspective, it is not surprising that the IDoS in hyperbolic lattices exhibits a quadratic behavior near the ground state, similar to the cubic lattice, see Fig. 6(b). We note that the DoS of hyperbolic lattices has been studied in recent literature. Mosseri et al. [106] investigate the DoS of {$p$,3} lattices by computing continous-fraction expansions of the lattice Green’s functions. They mention how a direct diagonalization technique (such as used in our work) results in a different DoS, due to boundary effects, which do not vanish in the thermodynamic limit in hyperbolic lattices. The fraction of sites that lie on the boundary of the graph goes to a constant as the size increases.
@@ III.4 Condensate fluctuations
Figure 7: Excited particles fluctuations in various geometries. Panel (a) shows the fluctuations in a cubic lattice with $V=15^{3}$ sites as a function of the temperature, peaking at the critical temperature $T_{c}$. Panel (b) shows the system size scaling scaling of fluctuations in the cubic geometry. For a set temperature equal to $1.5T_{c}$, the fluctuations scale linearly. For a temperature lower than $T_{c}$ the fluctuations are no longer linear, which is supported by the fitted function $b\cdot V^{a}$, with the resulting fitted parameter $a$ close to $4/3$, matches the continuum value in the thermodynamic limit [128]. In panels (c-j), we show the obtained fit parameter $a$, with error bars that reflect the statistical error of the fit, at various temperatures for: (c) cubic lattices, (d) square lattices, (e) Sierpiński triangular lattices, (f) regular triangular lattices, (g) Sierpiński tetrahedral lattices, (h) regular tetrahedral lattices, (i) $\{p,q\}=\{3,7\}$ hyperbolic lattices and (j) $\{p,q\}=\{7,3\}$ hyperbolic lattices. In regular lattices (panels c,d,f, and h), anomalous scaling is observed below $T_{\rm c}$, whereas above $T_{\rm c}$ the fluctuations scale normally. In fractal lattices (panels e and g), the exponent drops slowly, and fluctuations remain anomalous $(a>1)$ even for $T\gg T_{c}$. In the hyperbolic lattices (panels i and j), below the critical temperature the fitted function $\Delta N_{\text{ex}}(V)=bV^{a}$ does not describe correctly the excited particle fluctuations scaling. This is reflected by the large uncertainties of the fitted parameter. One can still observe a peak around the critical temperature, similarly to the tetrahedral lattice on panel (h).
Bose-Einstein condensates are not only characterized by the macroscopic occupation of the ground state, but also by anomalous fluctuations of the occupation numbers. The fluctuations of particles out of the condensate, $\Delta N_{\rm ex}^{2}$, are evaluated as [128]
@@
In the non-condensed regime, these fluctuations behave normally, that is, they scale linearly with the system size, $\Delta N_{\rm ex}^{2}\propto V$, where $V$ denotes the volume of the system, or the number of sites in the lattice. For a continuum gas in 3D, the anomalous fluctuations reflect in the behavior $\Delta N_{\rm ex}^{2}\propto V^{4/3}$, cf. [128]. As shown in Fig. 7(a-d), a similar anomalous scaling of fluctuations can also be observed in cubic lattices below $T_{c}$, and, with a different exponent, in square lattices.
In order to analyze fluctuations in the different lattice geometries, we have fitted the fluctuations towards a function $bV^{a}$, with $a,b$ fit parameters. The value of the exponent $a$ is plotted vs. temperature in Fig. 7(c-j) for regular, fractal and hyperbolic lattices.
Both for regular Euclidean geometries (Fig. 7(c,d,f,h)) and for hyperbolic lattices (Fig. 7(c,d,f,h)(i,j)), we find that the normal scaling regime, characterized by $a=1$, is relatively abruptly acquired as the temperature exceeds $T_{\rm c}$. On the other hand, in the fractal geometries (Fig. 7(e,g)), the normal regime is only slowly approached for $T\gg T_{\rm c}$. This behavior is in line with the heavy tail of the condensate fraction in fractal lattices.
Both for the fractal and the regular lattices, we observe strong deviations from normal scaling at low temperatures: for the tetrahedron grid, we have $a\approx 4/3$, in line with the expectation for a 3D gas. In a triangular lattice, the anomalous fluctuations are more pronounced, $a\approx 2$, and in fractal lattices, we even have $a>2$ at sufficiently small $T$. This is quite different from the behavior in hyperbolic lattices, where $a\approx 1$ (in the $\{p,q\}=\{3,7\}$ lattice) and $a\approx 1.1$ (in the $\{p,q\}=\{7,3\}$ lattice) is observed. Although these values are close (or even equal) to the value of normal scaling, the distinction between condensate phase and non-condensed phase is possible also in hyperbolic lattices, as $a$ exhibits a pronounced peak at $T_{\rm c}$.
@@ III.5 Long-range order
Figure 8: Thermal average of the single-particle density matrix (SPDM) elements in a Sierpiński triangle lattice with $V=366$ sites and a regular triangular lattice with $V=378$ sites. We plot $\rho_{ij}$, normalized by $\max_{i\neq j}|\rho_{ij}|$, for a fixed $j$, taken to be the site highlighted in red. In each column, two geometries are compared at the same absolute temperature. In the left column, at $T=0.075J/k_{B}$, we are far below $T_{\rm c}$ for both fractal and regular geometry, and both structures show long-range order. In the middle column, at $T=0.75J/k_{B}$, the fractal lattice is at criticality, and the long-range order in the fractal is reduced to a part of the fractal. In the right panel. at $T=2.3J/k_{B}=0.5T_{c-\text{triangle}}$, long-range order gets reduced also in the regular lattice.
[…]
As is shown in Fig. 8, long-range order is not only observed in regular lattices, but also for a condensate in the fractal lattice. In accordance with the lower critical temperature in the fractal lattice, the long-range order of the fractal is less robust against temperature as compared to the regular lattice.
@@ III.6 Discussion
We have compared the condensation behavior of ideal bosons in different lattice geometries (Euclidean lattices, fractal lattices, hyperbolic lattices). In all fractal lattices studied (Sierpiński gasket, Sierpiński carpet, Sierpiński tetrahedron), we find a significant suppression of critical temperature as compared to regular lattice systems. However, it should be noted that the condensate fraction exhibits a heavy tail, and a significant portion of the particles remains in the ground state even above $T_{\rm c}$. This behavior is a result of fragmentation of the condensate among several states within a quasi-degenerate manifold, as seen explicitly in the case of the Sierpiński carpet, cf. Fig. 3(c). In the thermodynamic limit (where no finite-size tail should appear), we do not expect condensation in the fractal lattice at finite temperature (see Fig. 4(a,c)), in accordance with their low dimensionality. This is also the case for the Sierpiński tetrahedron, which is embedded in a 3D space, but with a fractal dimension of 2. More specifically, the absence of condensation in the thermodynamic limit is a consequence of the linear scaling of integrated density of states at low energies, see Fig. 6(a). The condensates in the fractal are characterized by highly anomalous fluctuations of condensate occupation number, $\Delta N^{2}\sim V^{a}$ with $a>2.5$, see Fig. 7(e,g). Similarly to condensates in regular geometries, the fractal condensates exhibit off-diagonal long-range order in the one-body density matrix, see Fig. 8.
@@ IV.1 Cluster Gutzwiller ansatz
Let us start with the definition of a cluster. The cluster ${\cal C}$ on a lattice ${\cal L}$ is defined as a set of nodes forming a unit cell that can tile the whole lattice while preserving the structure of the node connections, for example see Fig. 9. As such, we define a cluster as a pair ${\cal C}=({\cal V_{\cal C}},{\cal E}_{\cal C})$ containing set of nodes ${\cal V}_{\cal C}=\{i,i=1,\dots,V_{\cal C}\}$, and edges ${\cal E}_{\cal C}=\{\langle i,j\rangle,i,j\in{\cal V}_{\cal C}\}$, with corresponding adjacency matrix $J^{\cal C}_{ij}=1$ for any $\langle i,j\rangle\in{\cal E}$, and 00 otherwise. We define the Bose-Hubbard Hamiltonian on the cluster ${\cal{C}}$
$\hat{H}_{\cal C}=$ $-J\sum_{i,j\in{\cal V}_{\cal C}}J^{\cal C}_{ij}(\hat{b}_{i}^{\dagger}\hat{b}_{j}+\hat{b}_{j}^{\dagger}\hat{b}_{i})+\frac{U}{2}\sum_{i\in{\cal V}_{\cal C}}\hat{n}_{i}(\hat{n}_{i}-1)$
@@
with bosonic Hilbert space ${\cal H}=\bigoplus_{i}{\cal H}_{N_{i},V_{\cal C}}$, where ${\cal H}_{N_{i},V_{\cal C}}$ is a Hilbert space of $N_{i}$ bosons on $V_{\cal C}$ sites. In this work, we consider $N/V_{\cal C}=1$ and $N_{i}=N-5,\dots,N,\dots,N+5$, and truncate the basis, by discarding Fock states with large local or global fluctuations, where $\exists_{i}\langle\psi||\hat{n}_{i}-V_{\cal C}/L||\psi\rangle>3$ or $\langle\psi|\sum_{i}|\hat{n}_{i}-V_{\cal C}/L||\psi\rangle>5$. Next, we define Hamiltonian describing the hopping between clusters, assuming the mean-field decoupling $\hat{b}_{i}^{\dagger}\hat{b}_{j}\approx\langle\hat{b}_{j}^{\dagger}\rangle\hat{b}_{i}\equiv\Phi^{*}\hat{b}_{i}$,
@@
where $\cal{V}_{\partial\cal C}$ is the set of indices belonging to the boundary of cluster ${\cal C}$ (denoted as $\partial\cal C$), $\nu_{i}$ is the number of external couplings of site $i$ to the mean-field (denoted as green edges on Fig 10 and Fig 11), and $\Phi=\langle GS|\hat{b}_{i_{0}}|GS\rangle$, where $i_{0}$ is an index of a site in the center of the cluster, and $|GS\rangle$ is a ground state of a coupled-cluster Hamiltonian
^$\hat{\cal{H}}_{{\cal C}}=\hat{H}_{\cal C}+\hat{H}_{\partial\cal C},$ (9)
The mean-field order parameter $\Phi$ is a signature of quantum phase of the system. It vanishes for Mott Insulator, and is non-zero for superfluid phase. The order parameter $\Phi$ can be obtained self-consistently starting with initial small value (here we initialize it as $\Phi=10^{-3}$), and iteratively updated until convergence. Since $\Phi$ converges monotonically (cf. [118]), we can distinguish the phase after a single iteration.
Figure 10: Phase diagrams of the Mott insulator and superfluid phases on $\{p,q\}=\{7,3\}$ (a) and $\{p,q\}=\{3,7\}$ (b) hyperbolic lattices for clusters (with $V_{\cal C}=16$, and $V_{\cal C}=15$ sites, respectively) shown above each plot. Green edges indicate mean-field couplings to the infinite lattice. The black regions denote the MI phase (vanishing order parameter $|\Phi|=0$), while yellow marks the SF phase where (non-zero order parameter $|\Phi|$).
We start with hyperbolic lattices. Figure 10 presents MI/SF phase diagram for two types of hyperbolic lattices, namely $\{p,q\}=\{7,3\}$ and $\{p,q\}=\{3,7\}$ on clusters with $V_{\cal C}=16$ and $15$ sites respectively, where $p$ denotes number of connections to the node, and $q$ denotes the number of edges in a lattice tile. To enable comparison across different geometries, the hopping amplitude is rescaled by the average coordination number $z$, which accounts for the total number of connections per site. The shapes of the lobes are similar to those in 2D Euclidean geometry. The $\{p,q\}=\{7,3\}$ lattice has a slightly taller lobe, which can be attributed to the smaller ratio of mean-field couplings to the total number of connections $\lambda$, as discussed in [120].
@@
Fig. 11, presents corresponding phase diagrams, allowing to see the quantitative change of the Mott lobe for three different Hausdorf dimensions $d=1,\log 3/\log 2,2$. Each cluster consists of $V_{\cal C}=15$ sites and is embedded in the surrounding lattice via the mean-field links, highlighted in green in the diagrams above the plots. Despite differences in internal structure, all sites in each cluster have the same total number of connections, which is $z=2$ in 1D, $z=4$ in the Sierpiński triangle, and $z=6$ in the triangular lattice. The rounded shape of the lobe in 2D and the sharper, more pointed shape in 1D are consistent with previous results [122, 123, 124]. The lobe in the Sierpiński lattice stands out by exhibiting features lying between the 1D and 2D geometries, reflecting the lattice’s non-integer fractal dimensionality.
Figure 11: Same as Fig.10, for lattices with dimension: $d=1$ – 1D lattice (a), $d=\log 3/\log 2$ – Sierpiński triangle (b), $d=2$ – triangular lattice (c). The Mott lobe for the fractal lattice exhibits a shape that lies in-between the one- and two-dimensional geometries, in accordance with its fractional Hausdorff dimension. Each cluster has $V_{\cal C}=15$ sites.

### v2

@@ III.2 Critical temperature
From the discussion of the condensate fraction, we have already noted substantial differences of the critical temperature in different geometries at fixed system sizes. A systematic comparison of $T_{\rm c}$ as a function of system size $V$ is shown in Fig. 4 for different fractal lattices, as well as for a regular 1D lattice (line), regular 2D lattices (triangular and square lattice), and regular 3D lattices (cubic and tetrahedral lattices). Not surprisingly, the latter ones are the only ones in which $T_{\rm c}$ does not decrease as a function of $V$. In fact, it is well established that, in the absence of a trapping potential, the critical temperature of a 2D system vanishes in the thermodynamic limit, and remains finite only in 3 (or more) spatial dimensions. More remarkably, we also observe in Fig. 4 that $T_{\rm c}$ is strictly ordered according to the Hausdorff dimension of each structure, both regarding its absolute value at a given $V$ and regarding its slope as a function of $V$ – with the exception of the Sierpiński carpet. The critical temperature for the carpet lattice decreases almost as rapidly as in the 1D lattice, due to condensate fragmentation discussed in the previous section. When fragmentation is taken into account, the critical temperature falls between that of the Sierpiński triangle ($d\approx 1.585$) and the 2D lattices, with only slight overlap with the Sierpiński tetrahedron. We further note that the Hausdorff dimension does not uniquely determine the behavior of $T_{\rm c}$: the triangular, square, and Sierpiński tetrahedron lattices possess the same Hausdorff dimension $d=2$, but exhibit slightly different values of $T_{\rm c}$.
@@
Strikingly different is the behavior in hyperbolic lattices. Although these lattices can be mapped onto a 2D Poincaré disk, the critical temperatures are found to increase with the system size, for sufficiently large systems, see Fig. 5. Clearly, this behavior suggests that the critical temperature does not vanish in the thermodynamic limit. Instead, we expect that the critical temperature will saturate as in the cubic lattice, however at system sizes that are too large for our computations. In this sense, the hyperbolic lattices provide a remarkable exception to other 2D structures with vanishing $T_{\rm c}$. In the following subsection, we will show that the different condensation behavior is explained by the very different spectral behavior in fractal, hyperbolic, or regular lattices.
@@
Figure 6: Integrated density of states (IDoS) in (a) three fractal lattices: Sierpiński carpet with $V=4096$ sites (black), Sierpiński triangle with $V=3282$ sites (orange), and Sierpiński tetrahedron with $V=2050$ sites (purple), as well as (b) hyperbolic and regular rectangular lattices: cubic lattice with $V=20^{3}=8000$ sites (black), rectangular lattice with $V=90^{2}=8100$ sites (purple), $\{p,q\}=\{3,7\}$ hyperbolic lattice with $V=11173$ sites (orange) and $\{p,q\}=\{7,3\}$ hyperbolic lattices with $V=9136$ sites (blue). For each geometry, the energy of each state $E$ is normalized using the energy of the ground state $\epsilon_{0}$ and the highest energy $\epsilon^{*}=\max(\epsilon_{k})$, so that $\tilde{E}=(E-\epsilon_{0})/(\epsilon^{*}-\epsilon_{0})$. In fractal lattices, the IDoS increases rapidly near the ground state, as shown in the red frame of Fig. 6(a). As can be seen in Fig. 6(b), the IDoS in the rectangular lattice scales linearly near the ground state, while in the cubic and hyperbolic lattices it follows a power-law scaling $\rm{IDoS}(\tilde{E})\sim\tilde{E}^{\alpha}$ with $\alpha>1.5$.
@@ III.3 Density of states
For an analysis of spectral properties it is convenient to look at the integrated density of states (IDoS) in the different geometries. In Fig. 6(a), we examine different fractal lattices which share similar properties, including large gaps (i.e. horizontal lines in the IDos) and huge degenerate of quasi-degenerate manifolds (i.e. vertical lines/jumps in the IDos). We note that these features are less pronounced in the Sierpiński carpet. The IDoS of the cubic and square lattice as well as different hyperbolic lattices is shown in Fig. 6(b). Their behavior is much smoother, and at low energies can be well fitted by a power-law function for the cubic lattice and the hyperbolic lattices and a linear function for the square lattice. It is well established that a linear IDoS, corresponding to a non-zero density of states (DoS) near the ground state, prevents BEC in the thermodynamic limit [140, 141]. From this perspective, it is not surprising that the hyperbolic lattices exhibits very similar condensation behavior as the cubic lattice, with a condensate that remains stable in the thermodynamic limit. We note that the DoS of hyperbolic lattices has also been studied in recent literature – Mosseri et al. [110] investigate the DoS of {pp,3} lattices by computing continuous-fraction expansions of the lattice Green’s functions. A power-law fit of the integrated density of states, $\mathrm{IDoS}(\tilde{E})\sim\tilde{E}^{\alpha}$, to the lowest $5\%$ of the spectrum, shown in Fig. 6(b), yields $\alpha=1.93$ for the 3,7{3,7} hyperbolic lattice, $\alpha=1.66$ for the 7,3{7,3} hyperbolic lattice, and $\alpha=1.65$ for the cubic lattice. In the cubic lattice, the density of states scales approximately as the square root of the excitation energy [142], corresponding to an expected $\alpha=1.5$. The slightly higher fitted value results from fitting over a finite portion of the spectrum rather than the asymptotic low-energy regime.
Although the spectral gaps in fractal systems prevent from fitting their IDoS to smooth functions, one can nevertheless approximate their rough trend at low $E$ through a power-law fit. For all fractal structures analyzed in Fig. 6(a), the exponent is on the order of $1$, providing the same argument against condensation that holds in the case of the square lattice.
@@ III.4 Condensate fluctuations
Figure 7: Excited particles fluctuations in various geometries. Panel (a) shows the fluctuations in a cubic lattice with $V=15^{3}$ sites as a function of the temperature, peaking at the critical temperature $T_{\rm c}$. Panel (b) shows the system size scaling of fluctuations in the cubic geometry. For $T=1.5T_{\rm c}$, the fluctuations scale linearly, whereas for $T<T_{\rm c}$ the fluctuations scale anomalously. with an exponent $a\approx 4/3$. In panels (c-j), we show the obtained fit parameter $a$ with error bars from the fit at various temperatures for: (c) cubic lattices, (d) square lattices, (e) Sierpiński triangular lattices, (f) regular triangular lattices, (g) Sierpiński tetrahedral lattices, (h) regular tetrahedral lattices, (i) $\{p,q\}=\{3,7\}$ hyperbolic lattices and (j) $\{p,q\}=\{7,3\}$ hyperbolic lattices.
Bose-Einstein condensates are not only characterized by the macroscopic occupation of the ground state, but also by the fluctuations of the occupation numbers. The fluctuations of particles out of the condensate, $\Delta N_{\rm ex}^{2}$, are evaluated as [139]
@@
In the non-condensed regime, fluctuations of the ground state population behave normally, that is, they scale linearly with the system size, $\Delta N_{\rm ex}^{2}\propto V$, where $V$ denotes the volume of the system, or the number of sites in the lattice. In the condensed regime, fluctuations may become anomalous, subtly depending on geometric properties of the system, interactions, and even the choice of thermodynamic ensemble, see Ref. [39] for a review. For a non-interacting continuum gas in a 3D box potential, the anomalous fluctuations are characterized by an exponent $a=4/3$, that is by the behavior $\Delta N_{\rm ex}^{2}\propto V^{a}$, cf. [139].
Here, we numerically study fluctuations in different lattice geometries. Specifically, we have fitted the fluctuations towards a function $bV^{a}$, with $a,b$ fit parameters. The value of the exponent $a$ is plotted vs. temperature in Fig. 7 for regular, fractal and hyperbolic lattices. As shown in Fig. 7(a-c), the fluctuation scaling in a cubic lattice is similar to the 3D continuum system. Below $T_{\rm c}$, we observe an anomalous scaling of fluctuations with a relatively flat value of $a\approx 4/3$. With an even larger exponent $a\approx 2$, anomalous fluctuations are also observed in square lattices, see Fig. 7(d). The behavior in triangular lattices or tetrahedral lattices, shown in Figs. 7(f,h), is very similar to the square and cubic cases, respectively, suggesting that mostly the dimensionality of the lattice matters in case of regular lattices.
Significant differences can be observed for fractal and hyperbolic lattices:
  • Both for the Sierpiński triangle and the Sierpiński tetrahedron, Fig. 7(e,g), the anomalous scaling of fluctuations is significantly larger than in regular 2D lattices. At very low temperatures, the exponent $a$ reaches values $a>2.5$, and it remains larger than 2 at any $T<T_{\rm c}$. In contrast to the regular lattices, where the exponent $a$ is approximately flat, the exponent $a$ becomes strongly temperature-dependent in fractal lattices.
  • Also the behavior above $T_{\rm c}$ brings the fractal in contrast to the regular lattice: While in a regular lattice $a$ drops abruptly to 1 at $T\approx T_{\rm c}$, $a$ decreases only slowly in the fractal systems, and we have $a>1$ even for temperatures far above $T_{\rm c}$. This is suggestive of a crossover regime between a condensed and non-condensed phase, in line with the heavy tail of the condensate fraction.
  • The hyperbolic systems, shown in Fig. 7(i,j), exhibit no or only very weak anomalous fluctuation scaling. This makes their scaling behavior in condensed and non-condensed phases very similar, with $a\approx 1$.
  • The distinction between condensed and non-condensed regimes in the hyperbolic lattices can be obtained from the fact that $a$ exhibits a peak (at around $a\approx 1.15$) in the vicinity of the critical temperature. Such a peak is not unique to the hyperbolic lattices, and can also be observed in the case of a tetrahedral lattice.
@@ III.5 Long-range order
Figure 8: Thermal average of the single-particle density matrix (SPDM) elements in a Sierpiński triangle lattice with $V=366$ sites and a regular triangular lattice with $V=378$ sites. The plots show $\rho_{ij}$, normalized by $\max_{i\neq j}|\rho_{ij}|$, for a fixed reference site $j$, highlighted in green. Each column compares the two geometries at the same absolute temperature. In the left column, the temperature $T=0.075\,J/k_{B}$ is far below $T_{\rm c}$ for both fractal and regular geometry, and both structures exhibit long-range order. In the middle column, at $T=0.75\,J/k_{B}$, the fractal lattice is at its critical temperature, and the long-range order is confined to a part of the structure. In the right column, where $T=2.3\,J/k_{B}$ corresponds to half the critical temperature of the triangular lattice, long-range order is reduced in both the regular and fractal lattices.
All properties discussed so far rely exclusively on the energy spectrum of the tight-binding model in a given graph. However, as Bloch theorem does not apply to non-periodic lattice such as the fractal ones, also the eigenstates in such geometries can differ significantly from the eigenstates in regular lattices. In particular, fractal lattices also admit localized eigenstates. Here, we analyze whether this affects the spatial properties of the corresponding Bose-Einstein condensates, in particular, we study the single-particle density matrix (SPDM) $\rho_{ij}=\langle\hat{b}_{i}^{\dagger}\hat{b}_{j}\rangle$, where $\langle\cdot\rangle$ shall denote thermal averaging. For regular lattices, condensation into the zero-momentum mode establishes long-range order of the SPDM, i.e. $\rho_{ij}\rightarrow\rho_{0}>0$ for $|i-j|\rightarrow\infty$.
As is shown in Fig. 8, long-range order is not only observed in regular lattices, but also for a condensate in the fractal lattice. Each panel displays the thermal average of the SPDM elements $\rho_{ij}$, calculated with respect to a single reference site (highlighted in green). In this way, the plot illustrates how correlations decay with distance from this site across the lattice.
At low temperature (left column), both the regular and fractal lattices exhibit extended coherence, with correlations remaining large throughout the system, signaling the presence of long-range order. At intermediate temperature (middle column), equal to the critical temperature of the fractal lattice, correlations remain strong only within parts of the fractal structure, while in the regular lattice they remain unchanged. Finally, at high temperature (right column), long-range order disappears also in the regular lattice, consistent with thermal depletion of the condensate. In accordance with the lower critical temperature in the fractal lattice, the long-range order of the fractal is thus less robust against temperature as compared to the regular lattice.
@@ III.6 Discussion
We have compared the condensation behavior of ideal bosons in different lattice geometries (Euclidean lattices, fractal lattices, hyperbolic lattices). In all fractal lattices studied (Sierpiński gasket, Sierpiński carpet, Sierpiński tetrahedron), we find a significant suppression of critical temperature as compared to regular lattice systems. However, it should be noted that the condensate fraction exhibits a heavy tail, and a significant portion of the particles remains in the ground state even above $T_{\rm c}$. This behavior is a result of fragmentation of the condensate among several states within a quasi-degenerate manifold, as seen explicitly in the case of the Sierpiński carpet, cf. Fig. 3(c). In the thermodynamic limit (where no finite-size tail should appear), we do not expect condensation in the fractal lattice at finite temperature (see Fig. 4(a,c)), in accordance with their low dimensionality. This is also the case for the Sierpiński tetrahedron, which is embedded in a 3D space, but with a fractal dimension of 2. More specifically, the absence of condensation in the thermodynamic limit is a consequence of the linear scaling of integrated density of states at low energies, see Fig. 6(a). The condensates in the fractal are characterized by highly anomalous fluctuations of condensate occupation number, $\Delta N_{\rm ex}^{2}\sim V^{a}$ with $a>2.5$, see Fig. 7(e,g). Similarly to condensates in regular geometries, the fractal condensates exhibit off-diagonal long-range order in the one-body density matrix, see Fig. 8.
@@ IV.1 Cluster Gutzwiller ansatz
[…]
$\hat{H}_{\cal C}=$ $-J\sum_{i,j\in{\cal V}_{\cal C}}J^{\cal C}_{ij}\hat{b}_{i}^{\dagger}\hat{b}_{j}+\frac{U}{2}\sum_{i\in{\cal V}_{\cal C}}\hat{n}_{i}(\hat{n}_{i}-1)$
@@
[…]
@@
where $\cal{V}_{\partial\cal C}$ is the set of indices belonging to the boundary of cluster $\cal C$} (denoted as $\partial\cal C$), $\nu_{i}$ is the number of external couplings of site $i$ to the mean-field (denoted as green edges in Fig 10 and Fig 11), and $\phi=\langle GS|\hat{b}_{i_{0}}|GS\rangle$, where $i_{0}$ is an index of a site in the center of the cluster, and $|GS\rangle$ is a ground state of a coupled-cluster Hamiltonian
$\hat{\cal{H}}_{{\cal C}}=\hat{H}_{\cal C}+\hat{H}_{\partial\cal C},$ (9)
The mean-field order parameter $\phi$ is a signature of the quantum phase of the system. It vanishes for Mott Insulator, and is non-zero for superfluid phase. The order parameter $\phi$ can be obtained self-consistently starting with initial small value (here we initialize it as $\phi=10^{-3}$), and iteratively updated until convergence. Since $\phi$ converges monotonically (cf. [128]), we can distinguish the phase after a single iteration.
Figure 10: Phase diagrams of the Mott insulator and superfluid phases on $\{p,q\}=\{7,3\}$ (a) and $\{p,q\}=\{3,7\}$ (b) hyperbolic lattices for clusters (with $V_{\cal C}=16$, and $V_{\cal C}=15$ sites, respectively) shown above each plot. Green edges indicate mean-field couplings to the infinite lattice. The black regions denote the MI phase (vanishing order parameter $|\phi|=0$), while yellow marks the SF phase where (non-zero order parameter $|\phi|$). The magenta line represents the strong-coupling (SC) expansion from Ref. [143], calculated for the $\{7,3\}$ lattice with $V=9136$ sites and the $\{3,7\}$ lattice with $V=11173$ sites.
We start with hyperbolic lattices. Figure 10 presents the MI/SF phase diagram for two types of hyperbolic lattices, namely $\{p,q\}=\{7,3\}$ and $\{p,q\}=\{3,7\}$ on clusters with $V_{\cal C}=16$ and $15$ sites respectively, where $p$ denotes number of connections to the node, and $q$ denotes the number of edges in a lattice tile. To enable comparison across different geometries, the hopping amplitude is rescaled by the average coordination number $z$, which accounts for the total number of connections per site. The shapes of the lobes are similar to those in 2D Euclidean geometry. The $\{p,q\}=\{7,3\}$ lattice has a slightly taller lobe, which can be attributed to the smaller ratio of mean-field couplings to the total number of connections $\lambda$, as discussed in [130].
@@
Fig. 11, presents corresponding phase diagrams, allowing to see the quantitative change of the Mott lobe for three different Hausdorff dimensions $d=1,\log 3/\log 2,2$. Each cluster consists of $V_{\cal C}=15$ sites and is embedded in the surrounding lattice via the mean-field links, highlighted in green in the diagrams above the plots. Despite differences in internal structure, all sites in each cluster have the same total number of connections, which is $z=2$ in 1D, $z=4$ in the Sierpiński triangle, and $z=6$ in the triangular lattice. The rounded shape of the lobe in 2D and the sharper, more pointed shape in 1D are consistent with previous results [133, 134, 135]. The lobe in the Sierpiński lattice stands out by exhibiting features lying between the 1D and 2D geometries, reflecting the lattice’s non-integer fractal dimensionality.
The MI/SF phase boundary can be approximated using a strong-coupling expansion, as described in Ref. [143]. By solving Eqs. (9) and (10) following the method described in detail in that reference, one can obtain an approximate boundary for lattices of arbitrary geometry. In Figures 10 and 11, the solid line represents the phase boundary calculated using this method for larger lattices of the corresponding geometry. In Figure 11, we see that the strong-coupling method reproduces the results for 1D and 2D lattices in a manner consistent with Ref. [143], where it is more accurate for the 1D geometry. On the other hand, the cluster mean-field method becomes more accurate in higher dimensions, which leads to near-perfect agreement between the two methods for the Sierpiński triangle lattice shown in Figure 11(b). In Figure 10, the strong-coupling approach yields slightly different results for the two hyperbolic geometries, and in both cases deviates from the cluster mean-field phase diagram near the tip of the Mott lobe.
Figure 11: Same as Fig.10, for lattices with dimension: $d=1$ – 1D lattice (a), $d=\log 3/\log 2$ – Sierpiński triangle (b), $d=2$ – triangular lattice (c). The Mott lobe for the fractal lattice exhibits a shape that lies in-between the one- and two-dimensional geometries, in accordance with its fractional Hausdorff dimension. Each cluster has $V_{\cal C}=15$ sites. The magenta line represents the strong-coupling (SC) expansion from Ref. [143], calculated for the 1D, Sierpiński triangle, and triangular lattices with $V=3282$, $V=3282$, and $V=3321$ sites, respectively.